OpenStudy (anonymous):

Differentiate f (x) f(x) = 3/x^2 + x + 4

OpenStudy (anonymous):

{f(x)=\frac{3}{x^2}+3} \cr{f(x)={3}{x^{-2}}+3}\cr

OpenStudy (anonymous):

\[{f(x)=\frac{3}{x^2}+x+4}=>{f(x)={3}{x^{-2}}+x+4}=>{f'(x)={-6}{x^{-3}+{1}}}\]

OpenStudy (anonymous):

Last part got cut off: \[{f'(x)={-6}{x^{-3}+{1}}}\] (if I'm remembering my integration correctly.)

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

thank you !